# On the Convergence of Multi-objective Optimization under Generalized Smoothness

### v3

@@ Abstract
Multi-objective optimization (MOO) is receiving more attention in various fields such as multi-task learning. Recent works provide some effective algorithms with theoretical analysis but they are limited by the standard $L$-smooth or bounded-gradient assumptions, which are typically unsatisfactory for neural networks, such as Long short-term memory (LSTM) models and transformers. In this paper, we study a more general and realistic class of $\ell$-smooth loss functions, where $\ell$ is a general non-decreasing function of gradient norm. We develop two novel single-loop algorithms for $\ell$-smooth MOO problems, Generalized Smooth Multi-objective Gradient descent (GSMGrad) and its stochastic variant, Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad), which approximate the conflict-avoidant (CA) direction that maximizes the minimum improvement among objectives. We provide a comprehensive convergence analysis of both algorithms and show that they converge to an $\epsilon$-accurate Pareto stationary point with a guaranteed $\epsilon$-level average CA distance (i.e., the gap between the updating direction and the CA direction) over all iterations, where totally $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples are needed for deterministic and stochastic settings, respectively. Our algorithms can also guarantee a tighter $\epsilon$-level CA distance in each iteration using more samples. Moreover, we propose an efficient variant of GSMGrad named GSMGrad-FA using only constant-level time and space, while achieving the same performance guarantee as GSMGrad. Our experiments validate our theory and demonstrate the effectiveness of the proposed methods.
@@ 1 Introduction
However, all existing works are limited by the standard $L$-smooth and bounded-gradient assumptions. Nevertheless, a recent study [35] indicates that such assumptions may not necessarily be true for the training of neural networks and an alternative $(L_{0},L_{1})$-smoothness condition was observed and studied, which assumes the Lipschitz constant to be linear in the gradient norm and the gradient norm to be potentially infinite. The analysis of existing MOO methods cannot be generalized to this $(L_{0},L_{1})$-smoothness directly due to the possible unbounded smoothness or gradient norm. In addition, all existing works [29, 35, 22, 21, 19, 11, 9] in generalized smoothness are limited to the single task problems, which are fundamentally different from the MOO problems. All of the above methods can not be directly generalized to the MOO tasks studied in this paper, since even though each single task is generalized smooth, the linear combination of these tasks is not necessarily generalized smooth. In this paper, we aim to fill this gap by proposing novel MOO algorithms, which not only converge under the generalized smoothness condition but also mitigate gradient conflict effectively with a guaranteed sufficiently small CA distance.
@@ 1.1 Our Contributions
We propose two single-loop MOO methods, Generalized Smooth Multi-objective Gradient descent (GSMGrad) and its stochastic variant SGSMGrad, and provide them with a comprehensive convergence analysis under the generalized smoothness condition in different settings. Our detailed contributions are listed below.
Weakest assumptions in MOO. In this paper, we investigate the $\ell$-smooth assumption, where $\ell$ is a general non-decreasing function of gradient norm, and includes both the standard $L$-smooth and $(L_{0},L_{1})$-smooth assumptions as special cases. This assumption finds many applications, such as LSTM models [35], transformers [11], distributionally robust optimization [19] and higher-order polynomial functions [9]. In addition, we do not make any bounded-gradient assumption, which is required in previous analysis to ensure the bounded multi-gradient approximation. To the best of our knowledge, this is the first work to investigate generalized smoothness in MOO problems.
New single-loop algorithms. Both GSMGrad and SGSMGrad are easy to implement by updating the weights $w$ of objectives and model parameters $x$ simultaneously via a single-loop structure. A warm-start initialization sub-procedure is also introduced at the beginning of both algorithms to ensure a sufficiently small iteration-wise CA distance under the single-loop structure. We also propose a computation- and memory-efficient variant of GSMGrad named GSMGrad-FA by updating the objective weights $w$ using only forward passes of $F(\cdot)$ rather than gradient $\nabla F$, which effectively reduces $O(K)$ time and space to $O(1)$ without hurting the performance guarantee.
Convergence analysis and optimal complexity. We provide a comprehensive analysis of our proposed algorithms under the generalized $\ell$-smooth condition in both deterministic and stochastic settings. To achieve an $\epsilon$-accurate Pareto stationary point and an $\epsilon$-level average CA distance, we show that GSMGrad and SGSMGrad require $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples in the deterministic and stochastic settings, respectively. Furthermore, to achieve a more aggressive $\epsilon$-level iteration-wise CA distance, GSMGrad and SGSMGrad require an increased number of samples, on the order of $\mathcal{O}(\epsilon^{-11})$ and $\mathcal{O}(\epsilon^{-17})$ respectively, in both deterministic and stochastic scenarios due to smaller step sizes and mini-batch data sampling. Typically, achieving an $\epsilon$-level iteration-wise CA distance results in much higher sample complexity, such as $\mathcal{O}(\epsilon^{-24})$ in [13], $\mathcal{O}(\epsilon^{-16})$ in [6] and $\mathcal{O}(\epsilon^{-12})$ in [32] for non-convex stochastic setting. Moreover, we show that GSMGrad-FA achieves the same performance guarantee as GSMGrad.
Supportive experiments. Our experiments on the MTL benchmark Cityscapes [10] validate our theory and demonstrate the effectiveness of our proposed algorithms.
@@ 1.2 Related Works
Gradient-based multi-objective optimization. A variety of gradient manipulation techniques have emerged for simultaneous learning of multiple tasks. One prevalent category of methods adjusts the weights of various objectives according to factors such as uncertainty [20], gradient norm [7], and training complexity [17]. Methods based on MOO have garnered increased attention due to their systematic designs, enhanced training stability and model-agnostic nature. For instance, [30] framed Multi-Task Learning (MTL) as a MOO problem and introduced an optimization method akin to MGDA [12]. Afterward, many MGDA-based methods have been proposed to mitigate gradient conflict with promising empirical performance. Among them, PCGrad [34] avoids conflict by projecting the gradient of each task on the norm plane of other tasks. GradDrop [8] randomly drops out conflicted gradients. CAGrad [24] adds a constraint on the update direction to be close to the average gradient. NashMTL [27] and FairGrad [3] formulated MTL as a bargaining game and a resource allocation problem, respectively. Theoretically, [13] proposed a provably convergent stochastic MOO method named MoCo based on an auxiliary tracking variable for gradient approximation. [6] characterized the trade-off among optimization, generalization, and conflict avoidance in MOO. [32] proposed a stochastic MOO method named SDMGrad with a preference-oriented regularizer, and analyzed its convergence. However, all these works rely on the $L$-smoothness and bounded-gradient assumptions. The details can be founded in Table 1. This paper focuses on the MOO problems with generalized $\ell$-smooth objectives.
Method Smoothness 11footnotemark: 1 Assumption22footnotemark: 2 Sample Complexity
SMG [25] (LS) (BG) N/A33footnotemark: 3
CR-MOGM[36] (LS) (BF), (BG) $\mathcal{O}(\epsilon^{-4})$
MoCo[13] (LS) (BF), (BG) $\mathcal{O}(\epsilon^{-4})$
MoDo[6] (LS) (BG) $\mathcal{O}(\epsilon^{-4})$
SDMGrad[32] (LS) (BG) $\mathcal{O}(\epsilon^{-4})$
SGSMGrad (this paper) (GS) N/A $\mathcal{O}(\epsilon^{-4})$
Table 1: Comparison for existing stochastic methods for MOO problems to obtain an $\epsilon$–accurate Pareto stationary point. Explanation on the upper footmarks: $1:$ (LS) indicates that the objectives are standard $L$-smooth while (GS) the objectives are generalized $\ell$-smooth as defined in Definition 1; $2:$ (BF) shows that the bounded function value assumption is required and (BG) shows that the bounded gradient assumption is required; $3:$ the analysis in [25] focuses on convex objective functions.
@@
Generalized smoothness. The generalized $(L_{0},L_{1})$-smoothness was firstly proposed by [35], which was observed from extensive empirical experiments in training neural networks. A clipping algorithm was developed by[35] and the convergence rate was provided. Later, [19] analyzed the convergence of a normalized momentum method. The SPIDER algorithm was also applied to solve generalized smooth problems in [29, 9], where [9] studied a new notion of $\alpha$-symmetric generalized smoothness, which includes $(L_{0},L_{1})$-smoothness as a special case. Very recently, a new $\ell$-smoothness condition was studied in [21, 22], which is the weakest smoothness condition and includes all the smoothness conditions discussed above. However, all the existing works on generalized smoothness are limited to single-task optimizations and the understanding of MOO is insufficient. This paper provides the first study of MOO under the generalized $\ell$-smoothness condition.
@@ 2.1 Generalized smoothness
($\ell$-smoothness, Definition 1 in [21]). A real-valued differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $\ell$-smooth if $\|\nabla^{2}f(x)\|\leq\ell(\|\nabla f(x)\|)$ almost everywhere in $\mathcal{X}$, where :[0,+)($\ell:[0,+\infty)\rightarrow(0,+\infty)$ is a continuous non-decreasing function.
The $(L_{0},L_{1})$-smoothness is a special case of $\ell$-smoothness, where $\ell(a)=L_{0}+L_{1}a$. There is another definition of generalized smooth, which is widely used and is equivalent to the $\ell$-smoothness:
@@
In $B(x,r(\|\nabla f(x)\|))$, the function $f$ is also $L$-smooth where $L=\ell(\|\nabla f(x)\|)$. Proposition 3.2 in [21] shows that Definition 1 and Definition 2 are equivalent: An $(r,\ell)$-smooth function is $\ell$-smooth; and an $\ell$-smooth function satisfying Assumption 1 is $(r,m)$-smooth where $m(u):=\ell(u+a)$ and $r(u):=a/m(u)$ for any $a>0$.
@@ 2.2 Pareto concepts in multi-objective optimization (MOO)
2.3 Multiple-gradient descent algorithm (MGDA) and its stochastic variants
Deterministic MGDA. One of the big challenges of MOO is the gradient conflict, i.e., the gradients of different objectives may vary heavily in scale such that the largest gradient dominates the update direction. As a result, the performance of those objectives with smaller gradients [34] may be significantly compromised. Towards this end, we tend to find a balanced update direction for all objectives. Thus, we consider the minimum improvement across all objectives and maximize it by solving the following problem
@@ 2.3 Multiple-gradient descent algorithm (MGDA) and its stochastic variants
where $\mathcal{W}$ is the probability simplex over $[K]$, and the regularization term $-\frac{1}{2}\|d\|^{2}$ is to regulate the magnitude of our update direction. The solution to the problem in eq. 3 can be obtained by solving the following problem [32]
@@
The above approach has been widely used in e.g., deterministic MGDA and its variants such as CAGrad, and PCGrad [12, 34, 24].
Stochastic MGDA. SMG [25] is the first stochastic MGDA. It directly replaces the gradients with stochastic gradients and the update rule becomes
@@
All the works mentioned above require bounded gradients such as [6, 13, 32] or $L$-smoothness such as [3, 24, 27, 33]. Their analyses do not apply to $\ell$-smoothness objectives studied in this paper, since the Lipschitz constant is potentially infinity.
@@ 2.4 Conflict-avoidant (CA) direction and CA distance
The larger the CA distance is, the further the estimated update direction will be away from the CA direction, and the more conflict there will be. In single-loop algorithms, MoCo and MoDo [13, 6], the average CA distance over iteration is of the order of $\epsilon$ while the double-loop algorithm SDMGrad [32] guarantees an $\epsilon$-order CA distance in every iteration. In our work, we analyze the CA distance in both cases and provide convergence results accordingly.
3 Single-loop Algorithms for MOO Under Generalized Smoothness
In this section, we present our main algorithms: Generalized Smooth Multi-objective Gradient descent (GSMGrad) and Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad), both are easy to implement with a simple single-loop structure. We also introduce an efficient variant of GSMGrad with constant-level computational and memory costs.
3.1 Generalized Smooth Multi-objective Gradient descent (GSMGrad)
We start to adopt MGDA in our method by computing an approximated weight $w_{t}$ and an update direction $d_{t}$ according to eq. 4 where $t$ is the iteration number. However, since the optimal weight $w^{*}$ of the convex function is not unique, we deal with this issue by adding an $\ell_{2}$ regularization term and the problem becomes
@@ 3.1 Generalized Smooth Multi-objective Gradient descent (GSMGrad)
Besides the benefit of a unique solution, adding an $\ell_{2}$ regularization term also makes $w_{\rho}^{*}(x)$ Lipschitz continuous [13]. Note that $w^{*}(x)$ may not be Lipschitz continuous because $\nabla F(x)^{\top}\nabla F(x)$ may not be positive definite. Nevertheless, the analysis of CA distance is difficult because $w^{*}$ may not be Lipschitz continuous. Thus, we will characterize the gap between $w^{*}$ and $w_{\rho}^{*}$ plus the change of $w_{\rho}^{*}$ after adding this $\ell_{2}$ regularization term. As a result, the update rules become Lines 4-5 in Algorithm 1. We first update $w_{t}$ by a projected gradient descent process and compute the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to update model parameters.
For our single-loop algorithm, CA distance is proportional to the term $\|w_{t}-w_{t,\rho}^{*}\|$, which decreases as the algorithm iterates with some error terms controlled by appropriately chosen small step sizes. If we initialize $w_{0}$ randomly, $\|w_{0}-w_{0,\rho}^{*}\|$ will be a constant order, and so will the first CA distance. Meanwhile, we can only get an $\epsilon$-order CA distance after a certain iteration number $t^{\prime}>1$ when $\|w_{t^{\prime}}-w_{t^{\prime},\rho}^{*}\|$ takes an $\epsilon$ order. Thus, we add an extra warm start process using Algorithm 2 to guarantee the new $w_{0}$ is close enough to $w_{0,\rho}^{*}$ and a small level CA distance in every iteration. However, this warm start process is not needed if we only require a small averaged CA distance.
Algorithm 1 Generalized Smooth Multi-objective Gradient descent (GSMGrad)
@@
2:  $w_{0}$=Warm-start($w_{0}$, $x_{0}$, $\rho$) for small-level iteration-wise CA distance only
3:  for $t=0,1,...,T-1$ do
4:     $w_{t+1}=\Pi_{\mathcal{W}}\big{(}w_{t}-\beta[\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\rho w_{t}]\big{)}$
5:     $x_{t+1}=x_{t}-\alpha\nabla F(x_{t})w_{t}$
6:  end for
@@
3.2 Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad)
Under the stochastic setting, our algorithm keeps the same structure, having a warm start process and an update loop if we aim to control the CA distance in every iteration. In Algorithm 2, we do the same projected gradient descent without using stochastic gradients. This is because we only need to compute $\nabla F(x_{0})^{\top}\nabla F(x_{0})$ once and reuse it in the whole loop, which does not bring a computational burden. Then in the update loop, we update the weight and model parameters accordingly. We use a double-sampling strategy here to make the weight gradient estimator unbiased [32] such that $d_{t}$ is a near-unbiased multi-gradient $\mathbb{E}[\nabla G_{2}(x_{t})^{\top}\nabla G_{3}(x_{t})w_{t}+\rho w_{t}]=\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\rho w_{t},$ where $\nabla G_{2}(x_{t})$ and $\nabla G_{3}(x_{t})$ are independent and unbiased estimates of $\nabla F(x_{t})$. Similarly, we do not involve a warm start process if we require the average CA distance to be small.
Algorithm 3 Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad)

### v4

# MGDA Converges under Generalized Smoothness, Provably

@@ Abstract
Multi-objective optimization (MOO) is receiving more attention in various fields such as multi-task learning. Recent works provide some effective algorithms with theoretical analysis but they are limited by the standard $L$-smooth or bounded-gradient assumptions, which typically do not hold for neural networks, such as Long short-term memory (LSTM) models and Transformers. In this paper, we study a more general and realistic class of generalized $\ell$-smooth loss functions, where $\ell$ is a general non-decreasing function of gradient norm. We revisit and analyze the fundamental multiple gradient descent algorithm (MGDA) and its stochastic version with double sampling for solving the generalized $\ell$-smooth MOO problems, which approximate the conflict-avoidant (CA) direction that maximizes the minimum improvement among objectives. We provide a comprehensive convergence analysis of these algorithms and show that they converge to an $\epsilon$-accurate Pareto stationary point with a guaranteed $\epsilon$-level average CA distance (i.e., the gap between the updating direction and the CA direction) over all iterations, where totally $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples are needed for deterministic and stochastic settings, respectively. We prove that they can also guarantee a tighter $\epsilon$-level CA distance in each iteration using more samples. Moreover, we analyze an efficient variant of MGDA named MGDA-FA using only $\mathcal{O}(1)$ time and space, while achieving the same performance guarantee as MGDA.
@@ 1 Introduction
Figure 1: Local smoothness constant vs. gradient norm for training SegNet on CityScapes dataset of Task 1.
However, all existing works are limited by the standard $L$-smooth and bounded-gradient assumptions. Nevertheless, a recent study [38] indicates that such assumptions may not necessarily be true for the training of neural networks and an alternative $(L_{0},L_{1})$-smoothness condition was observed and studied, which assumes the Lipschitz constant to be linear in the gradient norm and the gradient norm to be potentially infinite. Furthermore, this phenomenon has been consistently observed in our experiments (e.g., see Figure 1). Interestingly, it has been widely observed that MGDA algorithms always converge even under such generalized smoothness conditions [32, 25, 35]. This naturally raises a thought-provoking question:
  • Q.
    Can the fundamental MGDA algorithm provably converge under the generalized smoothness condition, while achieving a sufficiently small CA distance?
This question remains open due to the following challenges. The analysis of existing MOO methods cannot be generalized to this $(L_{0},L_{1})$-smoothness directly due to the possible unbounded smoothness or gradient norm. In addition, all existing works [31, 38, 22, 21, 19, 11, 9, 39] in generalized smoothness are limited to the single task problems, which are fundamentally different from the MOO problems since even though each single task is generalized smooth, the linear combination of these tasks is not necessarily generalized smooth. In this paper, we provide an affirmative answer to this question. Our main contributions are summarized below.
@@ 1.1 Our Contributions
We establish a comprehensive convergence analysis for MGDA under the generalized smoothness condition, in both the deterministic and stochastic settings. Moreover, the analysis covers both the average CA distance and the iteration-wise CA distance.
Weakest assumptions in MOO. In this paper, we investigate the generalized $\ell$-smooth assumption, where $\ell$ is a general non-decreasing function of gradient norm, and includes both the standard $L$-smooth and $(L_{0},L_{1})$-smooth assumptions as special cases. This assumption finds many applications, such as LSTM models [38], transformers [11], distributionally robust optimization [19] and higher-order polynomial functions [9]. In addition, we do not make any bounded-gradient assumption, which is required in previous analysis to ensure the bounded multi-gradient approximation. To the best of our knowledge, this is the first work to investigate generalized smoothness in MOO problems. Table 1 presents a detailed comparison of assumptions with existing analyses.
Convergence analysis. We first show that the vanilla MGDA method can converge to an $\epsilon$-accurate Pareto stationary point, while guaranteeing a small $\epsilon$-level average CA distance. A warm start process is introduced before the main loop of MGDA to achieve a more aggressive $\epsilon$-level iteration-wise CA distance. In the stochastic setting, we provide the convergence guarantee for MGDA with a double sampling, which was introduced by [35, 6] to obtain unbiased multi-gradient approximations. Furthermore, we analyze a computation- and memory-efficient variant of MGDA, named MGDA with fast approximation (MGDA-FA), which updates the objective weights $w$ using only forward passes of $F(\cdot)$ rather than gradient $\nabla F$, effectively reducing $O(K)$ time and space to $O(1)$ without hurting the performance guarantee.
Sample complexity comparison. To achieve an $\epsilon$-accurate Pareto stationary point and an $\epsilon$-level average CA distance, we show that MGDA require $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples in the deterministic and stochastic settings, respectively. Both of these complexities match with the existing best results. Furthermore, to achieve an $\epsilon$-level iteration-wise CA distance, MGDA with warm start requires an increased number of samples, at the order of $\mathcal{O}(\epsilon^{-11})$ and $\mathcal{O}(\epsilon^{-17})$ respectively, in both deterministic and stochastic scenarios, due to smaller step sizes and mini-batch data sampling. Typically, achieving an $\epsilon$-level iteration-wise CA distance results in much higher sample complexity, such as $\mathcal{O}(\epsilon^{-24})$ in [13], $\mathcal{O}(\epsilon^{-16})$ in [6] and $\mathcal{O}(\epsilon^{-12})$ in [35] for non-convex stochastic setting. Moreover, we show that MGDA-FA achieves the same performance guarantee as vanilla MGDA.
@@ 1.2 Related Works
Gradient-based multi-objective optimization. A variety of gradient manipulation techniques have emerged for simultaneous learning of multiple tasks. One prevalent category of methods adjusts the weights of various objectives according to factors such as uncertainty [20], gradient norm [7], and training complexity [17]. Methods based on MOO have garnered increased attention due to their systematic designs, enhanced training stability, and model-agnostic nature. For instance, [32] framed Multi-Task Learning (MTL) as a MOO problem and introduced an optimization method akin to MGDA [12]. Afterward, many MGDA-based methods have been proposed to mitigate gradient conflict with promising empirical performance. Among them, PCGrad [37] avoids conflict by projecting the gradient of each task on the norm plane of other tasks. GradDrop [8] randomly drops out conflicted gradients. CAGrad [25] adds a constraint on the update direction to be close to the average gradient. NashMTL [29] and FairGrad [3] formulated MTL as a bargaining game and a resource allocation problem, respectively. Theoretically, [13] proposed a provably convergent stochastic MOO method named MoCo based on an auxiliary tracking variable for gradient approximation. [6] characterized the trade-off among optimization, generalization, and conflict avoidance in MOO. [35] proposed and analyzed a stochastic MOO method named SDMGrad with a preference-oriented regularizer. However, all these works rely on the $L$-smoothness and bounded-gradient assumptions. In contrast, this paper focuses on the MOO problems with generalized $\ell$-smooth objectives.
Method Smoothness 11footnotemark: 1 Assumption22footnotemark: 2 Setting
CAGrad [25] (LS) (OD) Deterministic
PCGrad [37] (LS) (C) or (BC) Deterministic
SMG [27] (LS) (BG) Stochastic
CR-MOGM [40] (LS) (BF) and (BG) Stochastic
MoCo [13] (LS) (BF) and (BG) Stochastic
MoDo [6] (LS) (BG) Stochastic
SDMGrad [35] (LS) (BG) Stochastic
MGDA (Them. 1 and 3 ) (GS) N/A Deterministic
Stochastic MGDA (Them. 2 and 6) (GS) N/A Stochastic
[…]
@@
Generalized smoothness. The generalized $(L_{0},L_{1})$-smoothness was firstly proposed by [38], which was observed from extensive empirical experiments in training neural networks. A clipping algorithm was developed by [38] and the convergence rate was provided. Later, [19] analyzed the convergence of a normalized momentum method. The SPIDER algorithm was also applied to solve generalized smooth problems in [31, 9], where [9] studied a new notion of $\alpha$-symmetric generalized smoothness, which includes $(L_{0},L_{1})$-smoothness as a special case. Very recently, a new generalized $\ell$-smoothness condition was studied in [21, 22], which is the weakest smoothness condition and includes all the smoothness conditions discussed above. However, all the existing works on generalized smoothness are limited to single-task optimizations and the understanding of MOO is insufficient. This paper provides the first study of MOO under the generalized $\ell$-smoothness condition.
@@ 2.1 Generalized smoothness
(Generalized $\ell$-smoothness, Definition 1 in [21]). A real-valued differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$ is generalized $\ell$-smooth if $\|\nabla^{2}f(x)\|\leq\ell(\|\nabla f(x)\|)$ almost everywhere in $\mathcal{X}$, where :[0,+)($\ell:[0,+\infty)\rightarrow(0,+\infty)$ is a continuous non-decreasing function.
The $(L_{0},L_{1})$-smoothness is a special case of generalized $\ell$-smoothness, where $\ell(a)=L_{0}+L_{1}a$. There is another definition of generalized smooth, which is widely used and is equivalent to the $\ell$-smoothness:
@@
In $B(x,r(\|\nabla f(x)\|))$, the function $f$ is also $L$-smooth where $L=\ell(\|\nabla f(x)\|)$. Proposition 3.2 in [21] shows that Definition 1 and Definition 2 are equivalent: An $(r,\ell)$-smooth function is $\ell$-smooth; and a generalized $\ell$-smooth function satisfying Assumption 1 is $(r,m)$-smooth where $m(u):=\ell(u+a)$ and $r(u):=a/m(u)$ for any $a>0$.
@@ 2.2 Pareto concepts in multi-objective optimization (MOO)
2.3 Existing MOO Algorithms
Deterministic MOO. One of the big challenges of MOO is the gradient conflict, i.e., the gradients of different objectives may vary heavily in scale such that the largest gradient dominates the update direction. As a result, the performance of those objectives with smaller gradients [37] may be significantly compromised. As the most fundamental MOO algorithm, MGDA tends to find a balanced update direction for all objectives by considering the minimum improvement across all objectives and maximize it by solving the following problem
@@ 2.3 Existing MOO Algorithms
where $\mathcal{W}$ is the probability simplex over $[K]$, and the regularization term $-\frac{1}{2}\|d\|^{2}$ is to regulate the magnitude of our update direction. The solution to the problem in eq. 3 can be obtained by solving the following problem [12]:
@@
The above approach has been widely used in various variants of MGDA such as SDMGrad, CAGrad, and PCGrad [35, 37, 25].
Stochastic MOO. SMG [27] is the first stochastic MGDA. It directly replaces the gradients with stochastic gradients and the update rule becomes
@@
Strong assumptions in analysis. All the works mentioned above require bounded gradients such as [6, 13, 35] or $L$-smoothness such as [3, 25, 29, 36]. Their analyses do not apply to generalized $\ell$-smoothness objectives, since the Lipschitz constant is potentially infinity.
@@ 2.4 Conflict-avoidant (CA) direction and CA distance
The larger the CA distance is, the further the estimated update direction will be away from the CA direction, and the more conflict there will be. In single-loop algorithms, MoCo and MoDo [13, 6], the average CA distance over iteration is of the order of $\epsilon$ while the double-loop algorithm SDMGrad [35] guarantees an $\epsilon$-order CA distance in every iteration. In this work, we analyze the CA distance in both cases and provide convergence results.
3 MGDA Algorithms under Generalized Smoothness
3.1 MGDA with and without Warm Start
It has been shown in eq. 4 that MGDA needs to approximate the optimal weight $w^{*}$ and the optimal updating direction $d^{*}$. However, since the optimal weight $w^{*}$ of the convex function is not unique, we deal with this issue by adding an $\ell_{2}$ regularization term and the problem becomes
@@ 3.1 MGDA with and without Warm Start
Besides the benefit of a unique solution, adding an $\ell_{2}$ regularization term also makes $w_{\rho}^{*}(x)$ Lipschitz continuous [13]. Note that $w^{*}(x)$ may not be Lipschitz continuous because $\nabla F(x)^{\top}\nabla F(x)$ may not be positive definite. Nevertheless, the analysis of CA distance is difficult because $w^{*}$ may not be Lipschitz continuous. Thus, we will characterize the gap between $w^{*}$ and $w_{\rho}^{*}$ plus the change of $w_{\rho}^{*}$ after adding this $\ell_{2}$ regularization term. As a result, the update rules become Lines 5-6 in Algorithm 1. We first update $w_{t}$ by a projected gradient descent process and compute the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to update model parameters.
For our single-loop algorithm, CA distance is proportional to the term $\|w_{t}-w_{t,\rho}^{*}\|$, which decreases as the algorithm iterates with some error terms controlled by appropriately chosen small step sizes. If we initialize $w_{0}$ randomly, $\|w_{0}-w_{0,\rho}^{*}\|$ will be a constant order, and so will the first CA distance. Meanwhile, we can only get an $\epsilon$-order CA distance after a certain iteration number $t^{\prime}>1$ when $\|w_{t^{\prime}}-w_{t^{\prime},\rho}^{*}\|$ takes an $\epsilon$ order. Thus, we introduce an extra warm start process using Algorithm 2 to guarantee the new $w_{0}$ is close enough to $w_{0,\rho}^{*}$ and a small level CA distance in every iteration. However, this warm start process is not needed if we only require a small averaged CA distance.
Algorithm 1 Single loop MGDA with and without warm start
@@
2:  Option I: $w_{0}$=warm-start($w_{0}$, $x_{0}$, $\rho$) # for analyzing iteration-wise CA distance
3:  Option II: $w_{0}\leftarrow w_{0}$                      # for analyzing averaged CA distance
4:  for $t=0,1,...,T-1$ do
5:     $w_{t+1}=\Pi_{\mathcal{W}}\big{(}w_{t}-\beta[\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\rho w_{t}]\big{)}$
6:     $x_{t+1}=x_{t}-\alpha\nabla F(x_{t})w_{t}$
7:  end for
@@
3.2 Stochastic MGDA with Double Sampling
In the stochastic setting, our algorithm keeps the same structure, having a warm start process if we aim to control the CA distance in every iteration. In Algorithm 2, we do the same projected gradient descent without using stochastic gradients. This is because we only need to compute $\nabla F(x_{0})^{\top}\nabla F(x_{0})$ once and reuse it in the whole loop, which does not bring a computational burden. Then in the update loop, we update the weight and model parameters accordingly. We use a double-sampling strategy here to make the weight gradient estimator unbiased [35] such that $d_{t}$ is a near-unbiased multi-gradient $\mathbb{E}[\nabla G_{2}(x_{t})^{\top}\nabla G_{3}(x_{t})w_{t}+\rho w_{t}]=\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\rho w_{t},$ where $\nabla G_{2}(x_{t})$ and $\nabla G_{3}(x_{t})$ are independent and unbiased estimates of $\nabla F(x_{t})$. Similarly, we do not involve a warm start process if we require the average CA distance to be small.
Algorithm 3 Stochastic MGDA with Double Sampling